[figure: 9.7 cm]

MAGNETIC PROPERTIES OF Cr_{3}Si DOPED F e AND Co

P. Zaleski^{a}, M. Biernacka^{a}, L. Dobrzynski^{a,b}, K. Perzynska^{a}, K. Recko^{a} aUniversity of Bialystok, Institute of Experimental Physics, 41 Lipowa str. Bialystok,

Poland bThe Soltan Institute of Nuclear Studies, Otwock, Swierk, Poland

Policrystalline samples of Cr_{3−x}T M_{x}Si (where x=0.02, 0.1, 0.2, 0.3; TM=Fe, Co) were prepared from pure elements in the arc furnace. The magnetization of these samples was measured by VSM method as a function of magnetic field (-12 kOe to 12 kOe) and temperature (10 K to 300 K). With increasing iron concentration the Cr_{3−x}F e_{x}Si alloys change from the paramagnetic (x≤0.2) to the ferromagnetic (x=0.3) ones. In the case of Co impurities the ferromagnetic properties were observed in the most diluted samples (x=0.02).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : P. Zaleski

Address for correspondence : University of Bialystok Institute of Experimental Physics 41 Lipowa str. 15-424 Bialystok, Poland

Email address : pzal@alpha.uwb.edu.pl